OpenStudy (anonymous):

The height of the tunnel at the center is 54 ft and the vertical clearance must be 18 ft at a point 8 ft from the center. Find an equation for the ellipse.

OpenStudy (goformit100):

Madam Can I Help you?

OpenStudy (anonymous):

You can! @goformit100

OpenStudy (nurali):

X^2/A^2 + Y^2/B^2 = 1 12^2/A^2 + (18/54)^2 = 1 144/A^2 = 1 - 1/9 = 8/9 A^2 = 9*144/8=162 Therefore the eq is X^2/162 + Y^2/2916 = 1

OpenStudy (anonymous):

THANK YOU @Nurali

OpenStudy (anonymous):

wait a second that does not look correct when graphed! @Nurali

OpenStudy (anonymous):

It must look like this! @Nurali

OpenStudy (anonymous):

Oh I'm sorry! Mathway made it look funny :( @Nurali

OpenStudy (nurali):

no problem.